### v2

# RGBD Objects in the Wild: Scaling Real-World 3D Object Learning from RGB-D Videos

Abstract
We introduce a new RGB-D object dataset captured in the wild called WildRGB-D. Unlike most existing real-world object-centric datasets which only come with RGB capturing, the direct capture of the depth channel allows better 3D annotations and broader downstream applications. WildRGB-D comprises large-scale category-level RGB-D object videos, which are taken using an iPhone to go around the objects in 360 degrees. It contains around 8500 recorded objects and nearly 20000 RGB-D videos across 46 common object categories. These videos are taken with diverse cluttered backgrounds with three setups to cover as many real-world scenarios as possible: (i) a single object in one video; (ii) multiple objects in one video; and (iii) an object with a static hand in one video. The dataset is annotated with object masks, real-world scale camera poses, and reconstructed aggregated point clouds from RGBD videos. We benchmark four tasks with WildRGB-D including novel view synthesis, camera pose estimation, object 6d pose estimation, and object surface reconstruction. Our experiments show that the large-scale capture of RGB-D objects provides a large potential to advance 3D object learning. Our project page is https://wildrgbd.github.io/.
Figure 1: WildRGB-D Dataset contains almost 8500 recorded objects and nearly 20000 RGBD videos in 46 common categories with corresponding object masks and 3D point clouds.
00footnotetext: Equal contribution.
Figure 2: The camera poses trajectory in WildRGB-D Dataset. We visualize the corresponding camera in each scene of our dataset, showing that our dataset is featured in 360 degree full and dense multi-view camera poses.
1 Introduction
The recent advancement of computer vision has been largely relying on the scaling of training data [26, 52]. The same success in data-driven approaches has been recently adopted to 3D object modeling with new large 3D object-centric dataset collection [39, 51, 25]. Most of the large datasets are synthetic 3D data [5, 22, 74, 14] and a mix of synthetic data and real-world object scans [16], given it is much less labor intensive for scaling by rendering from simulation. However, it remains a big challenge to apply the model trained in simulation data to the real world. This is not only because the synthetic data has less realistic texture and shape, but also due to it is very hard to model the cluttered background and the natural light comes with it in simulation.
[…]
In this paper, we propose to collect a new dataset that contains large-scale RGB-D object videos across diverse object categories and presented in the wild. Our dataset, namely WildRGB-D, covers 8500 tabletop objects across 44 categories in 20K videos. The videos are taken using iPhones to go around the objects in 360 degrees (see Figure 2 for visualization). Examples of the dataset are shown in Figure 1. There are three types of videos: (i) Single object video where there is only one object presented on the table; (ii) Multi-object video where there are multiple objects presented at the same time; and (iii) Hand-object video where there is a static human hand grasping the object. The collection of the WildRGB-D  dataset not only considers the cluttered background in the real world, but also the common scenarios where the objects are occluded by human hands.
We perform automatic annotations for WildRGB-D. With RGB-D capturing, we can apply the Simultaneous Localization and Mapping [19, 57] (SLAM) algorithms, and exploit the RGB images and depth information from the depth sensor of mobile phones to reconstruct the 3D camera poses in real-world scale and aggregated 3D point clouds. Additionally, center object segmentation masks can be attained by bounding-box detection using text prompt of object category in Grounding-DINO [40], segmentation using Segment-Anything [32] and mask tracking using XMem [11], which are largely integrated into [12, 76].
To exploit the potential of our dataset, we benchmark it in four downstream tracks:
(i) Novel view synthesis. We evaluate various algorithms based on NeRF [46] which is optimized in a single scene, and generalizable NeRF which is trained in a category-level. With the help of depth information when training NeRFs, we can achieve consistently improved results. This offers a new platform for evaluating view synthesis approaches using RGB or RGB-D data.
(ii) Camera pose estimation. We adopt different pose estimation approaches [84, 34] to evaluate their capability of estimating relative camera poses in a sparse setting. We validate their generalizable ability through training on a partial of all categories and testing on unseen ones. We observe remarkable generalization performance on unseen categories, which indicates our large-scale category-level dataset can serve as a training source for generalizable camera pose estimation.
(iii) Object surface reconstruction. We conduct object surface reconstruction in our dataset with RGB or RGB-D videos and object masks through Instant-NGP [48] and Neusfacto [83]. Results show that our depth information endow reconstruction with more accurate precision and SDF-based algorithm [83] performs better in this setting.
(iv) Object 6D pose estimation. We exploit the self-supervised algorithm in category-level object 6D pose estimation [85] with large-scale RGB-D images in our dataset and then evaluate the pre-trained model on the Wild6D [23] test set. We show our dataset can facilitate 6D pose estimation even without training labels, and we also study its generalization ability to the out-of-distribution test set.
2 Related Work
Dataset Real Multi-View Depth Src. 3D GT Video # Cats # Objs
ShapeNet [5] none CAD mesh none 55 51k
ModelNet [74] none CAD mesh none 40 12k
3D-Future [22] none CAD mesh none 34 16k
ABO [14] none CAD mesh none 63 8k
DTU [1] limited SfM mesh RGB N/A 124
CO3D [53] full SfM pcl RGB 50 19k
MVImgNet [82] limited SfM pcl RGB 238 220k
Objectron [2] limited SfM pcl RGB 9 15k
GSO [18] none scanner mesh none 17 1k
OmniObject3D [73] limited scanner mesh RGB 190 6k
Choi et al. [13] limited sensor mesh* RGBD 9 2k
Wild6D [23] limited sensor pcl RGBD 5 1.8k
Ours full sensor pcl RGBD 44 8.5k
Table 1: Comparison of WildRGB-D dataset with other 3D object dataset. Some datasets don’t provide video and we mark in “none”. Some only cover partial angles, which is marked in “limited”. Asterisk(*) means partial annotations. Depth Src. means where the depth information comes from, including CAD models, SfM COLMAP, scanner devices and depth sensor in iPhones. pcl is the abbreviation of point cloud.
Figure 3: Statistics of WildRGB-D Dataset list the total and per-category number of objects and different types of videos.
Figure 4: Distribution visualization of different kinds of Object 6D pose dataset and WildRGB-D  dataset. We observe obvious disparity between Wild6D and Our dataset. Visualizations of  [66, 68, 41, 23, 30] are from  [30].
3D Object Datasets
One representative kind of 3D object dataset is 3D synthetic dataset, like ShapeNet [5] and ModelNet40 [74], which consist of category-level objects. 3D-FUTURE [22] and ABO [14] datasets are typical of higher quality mesh with textures. [33] and [63] introduce real-world category-specific object datasets that mainly focus on birds and chairs respectively. DTU [1] and BlendedMVS [78] are datasets designed for multi-view stereo and lack category-level classification. Objectron [2] provides rich annotations but only partial videos are fully 360 degree covered. CO3D [53] is a large-scale category-level dataset that annotates camera poses and depths with SfM tool COLMAP [56], which doesn’t provide depth information in real-world scale. Pascal3D [75] contains real-world 3D objects with pose annotations and CAD models in limited categories. Datasets collected with specialized hardware (scanner, dome, etc.) like GSO [18] and OmniObject3D [73] have more accurate 3D geometry models. However, they don’t have RGBD wild object videos collected. In the aspect of RGBD object datasets, Wild6D [23] features RGBD image sequences and 6D pose annotations while lacking full 360 coverage and category types. Choi et al. [13] proposes RGBD object-centric datasets in 44 categories, but with limited camera annotations. As a comparison, our proposed WildRGB-D  dataset contains almost 8500 recorded objects and nearly 20000 RGBD videos recorded all 360 degrees in 46 common categories from well-known 2D datasets, all with real-world scale camera poses and object mask annotations as well as aggregated point clouds. We present the detailed comparison in Tab. 1.
Neural Radiance Field and Object Surface Reconstruction
Neural Radiance Field (NeRF) [46] is a kind of scene representation based on MLPs. It takes in sampled points along each ray and outputs the density and color of each point, which are then aggregated by volume rendering to synthesize views. [3, 4, 47, 65] propose new changes to the original NeRF to improve the visual quality and [7, 21, 48, 59, 38] advance the NeRF efficiency. In order to generalize the NeRF representations to other scenes,  [81, 6, 69, 28, 42] learn latent 3D representations and priors from a bunch of existed scenes to help synthesize views across different scenes. Derived from original NeRF, [67, 79, 83, 15, 49, 72] leverage Sign Distance Function (SDF) and represent the 3D scene by implicit surface, which has a more clear object boundary definition. WildRGB-D  dataset comprises various category-level objects and scenes on a large scale, which is suitable for novel view synthesis benchmarks and helps boost more mature reconstruction algorithms and generalizable 3D scene representations.
Camera Pose Estimation
Given dense image views, mature algorithms of SfM [56] and SLAM [19] can estimate camera poses well by computing visual matches [43], verifying through RANSAC [20] and optimizing via bundle adjustment [62]. However, in a sparse camera view setting, camera pose estimation remains a challenging task. Some approaches [60, 70] leverage RNN or adopt auto-regression [77] targeting at SLAM applications. For category-agnostic sparse view camera pose estimation,  [45, 54] adopt a direct regression approach.  [29] estimates 6D pose upon training on synthetic dataset. Energy-based method [84] estimates distributions over relative rotations and  [34] incorporates multi-view context to estimate camera 6D pose. Bundle adjustment gets learned after predictions in  [58] to refine the estimated poses. In WildRGB-D  dataset, with full 360-degree multi-view videos, the sparse view camera pose estimation setting is easily accessible, enabling our dataset to serve as a large-scale training database for these algorithms.
Object 6D Pose Estimation
In the setting of category-level 6D pose estimation, algorithms predict object poses in the same category and meet with various intra-class shapes.  [66] predicts 6D pose using Umeyama algorithm [64] with NOCS map estimation and  [61, 8, 23] follow up to learn more accurate NOCS representations. Other algorithms learn to estimate 6D pose through direct regression [9, 35], keypoint location estimations [37] and so on. Apart from supervised learning, self-supervision emerges due to the high cost of annotations. One approach [10, 27, 24, 80] is to adapt sim-to-real upon the pre-trained model on synthetic data. Another one [23, 44, 50] resorts to semi-supervised training. [85] proposes cycles across 2D-3D space learned correspondence, which enables training using only in-the-wild RGBD images without any annotations and is compatible with our dataset. With large-scale category-level RGBD wild object images for self-supervised learning, our dataset has the potential to boost future developments in this field.
3 The WildRGB-D Dataset
Figure 5: Point cloud reconstruction of objects in WildRGB-D Dataset. We reconstruct the aggregated point cloud of the scene by leveraging existed 3D annotations of camera poses and depth images.
3.1 Data Collections, Processing, and Annotation
Data Collections
In order to collect RGBD video on a large scale expediently and economically, we record with the help of iPhone front camera using Record3D  [^1] https://record3d.app/ App and rotate the camera around the object so that full 360-degree views of objects are captured with RGB images and the corresponding depth images. Camera rotating speed is controlled equably to ensure less blur in videos. We select 46 common categories from well-known 2D datasets [17, 36]. We record three videos for every selected object, which are composed of single-object video, multi-object video, and hand-object video. Every recorded video has been checked and some are left behind due to poor SLAM camera pose estimation. Details of WildRGB-D Dataset are listed in Fig. 3.
Generating Camera Poses and 3D Point Cloud
Our WildRGB-D dataset has 3D annotations including camera poses in real-world scale, scene point clouds, and central object masks. In order to attain real-world scale camera poses, applying Structure-from-Motion (SfM) framework COLMAP [56] will result in relatively inaccurate SfM 3D camera pose annotations and the discrepancy of scaling in poses compared to real-world scaling, since SfM only leverages RGB image sequence and doesn’t take corresponding depths into account. The mature RGBD Simultaneous Localization and Mapping [19, 57] (SLAM) algorithms have the capability of exploiting the RGB images and depth information from depth sensor of mobile phones to reconstruct the 3D camera poses in real-world scale, which enables us to simply project the depth images and gain aggregated 3D point clouds (see Figure 5). Then we manually check the quality of the aggregated 3D point cloud and exclude videos in which SLAM fails to get accurate camera poses. To increase the probability of getting correct SLAM results for each video, we adopt two kinds of SLAM algorithms including BAD SLAM [57] and SLAM implementation from Open3D [87], which increase our successful rate to over 90%.
Generating Central Object Masks
We perform central object mask segmentation through a series of methods. Instead of the classic PointRend [31] algorithm, we leverage the novel segmentation tool Segment-Anything (SAM) [32]. We attain the prompt for SAM using Grounding-DINO [40], which generates a bounding box for SAM according to the category text prompt. After attaining the mask segmentation of the first frame in the video, XMem [11] is applied to track the mask in the video. The masking pipeline is largely integrated into [12, 76].
3.2 Statistics and Distribution
In WildRGB-D dataset collections, we recorded 8500 objects and 3 videos for each one. After excluding those SLAM-failed videos, we have 8367 objects in 23049 videos in our dataset (maintaining rates are 99.3%/91.0%). The selected videos contain 33.1% single object videos, 63.0% multi-object videos, and 3.9% hand-object videos. Details of WildRGB-D  dataset are listed in Fig. 3.
4 Experiments
[…]
NeRF [46] 23.03/1.50 0.690/0.072 0.390/0.075 0.306/0.109
NeRF (w mask) 34.65/4.44 0.943/0.077 0.031/0.032 0.029/0.019
Mip-NeRF 360 [4] 23.84/1.60 0.762/0.063 0.280/0.067 0.185/0.068
Mip-NeRF 360 (w mask) 35.60/4.51 0.949/0.077 0.024/0.025 0.020/0.015
Instant-NGP [48] 23.67/2.07 0.745/0.063 0.257/0.070 0.366/0.105
Instant-NGP (w mask) 35.65/5.20 0.946/0.077 0.021/0.031 0.068/0.074
Table 2: Single-scene NVS results. Average of four metrics w and w/o masks across all training dataset are reported with their standard deviation (SD).
[…]
Pixel-NeRF [81] Easy 20.28/0.65 0.645/0.043 0.495/0.074 0.355/0.120
MVSNeRF [6] 19.95/1.00 0.663/0.036 0.351/0.066 0.370/0.100
IBRNet [69] 20.93/0.98 0.711/0.031 0.395/0.153 -
Pixel-NeRF [81] Middle 18.76/0.50 0.572/0.064 0.534/0.047 0.299/0.057
MVSNeRF [6] 18.75/0.74 0.601/0.069 0.363/0.036 0.345/0.102
IBRNet [69] 19.77/1.01 0.663/0.071 0.362/0.063 -
Pixel-NeRF [81] Hard 17.23/0.66 0.521/0.035 0.624/0.054 0.383/0.121
MVSNeRF [6] 17.13/0.89 0.564/0.043 0.425/0.045 0.502/0.260
IBRNet [69] 17.92/1.12 0.614/0.056 0.439/0.069 -
Table 3: Cross-scene NVS results. Average of four metrics across all categories in training dataset are reported. We report metrics of three difficulty level respectively. Entries marked in - are not provided.
[…]
Instant-NGP [48] 23.67/2.07 0.745/0.063 0.257/0.070 0.366/0.105
Instant-NGP (depth sup.) 24.60/2.13 0.759/0.062 0.239/0.066 0.108/0.057
Pixel-NeRF [81] 18.53/1.21 0.568/0.067 0.556/0.073 0.336/0.099
Pixel-NeRF (depth sup.) 19.10/1.21 0.605/0.060 0.499/0.064 0.147/0.087
MVSNeRF [6] 18.43/1.30 0.600/0.065 0.381/0.054 0.400/0.182
MVSNeRF (depth sup.) 18.44/1.29 0.600/0.065 0.381/0.054 0.397/0.186
Table 4: Depth Supervised NVS and depth estimation results. Average of four metrics w and w/o depth supervision across all training dataset are reported with their standard deviation (SD).
Target NeRF [46] Mip-NeRF 360 [4] Instant-NGP [48]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Figure 6: Novel view synthesis visualization of different kinds of NeRF methods: NeRF [46], Mip-NeRF 360 [4] and Instant-NGP [48].
4.1 Novel View Synthesis
In this section, we conduct multiple experiments towards methods concerning novel view synthesis (NVS) in the following three scenarios: 1) Single-Scene NVS, where we train NeRF-based methods [46, 4, 48] on a single scene with only RGB image sequence. 2) Cross-Scene NVS, where we learn category-level scene representations to generalize into other scenes with Generalizable NeRFs [81, 6, 69]. 3) Depth Supervised NVS, where we conduct NVS experiments with depth image priors in our dataset to study the potential that depth information will endow to NVS tasks.
Single-Scene NVS
We select ten scenes from each category and uniformly sample images as validation split. We choose NeRF [46], Mip-NeRF 360 [4] and Instant-NGP [48] for evaluations. Results are shown in Tab. 2. We report the average PSNR, SSIM [71], LPIPS [86] and rendering depths Mean Average Error (MAE) compared with our sensor-collected depths across all categories. We also report metrics only related to the NVS quality of central objects using object masks. Results show that Mip-NeRF 360 and Instant-NGP outperform original NeRF in terms of visual quality metrics. NeRF-based methods perform better when we only concern with the recovery of central objects under object masks. What’s more, Mip-NeRF 360 performs best in learning single-scene geometry due to its special designs of proposal MLP and distortion loss. Visualization of single-scene NVS quality can be found in Figure 6. In brief, our dataset offers various and extensive categories and scenes for in-depth NVS experiments.
Cross-Scene NVS
Apart from single-scene NeRF optimizations, we also evaluate Generalizable NeRFs: Pixel-NeRF [81], MVSNeRF [6] and IBRNet [69] in the cross-scene setting. For each category in our dataset, we select the same ten test scenes as single-scene NVS experiments and train in the remaining scenes of the same category to learn per-category latent representations. We divide the 46 categories into three difficulty levels and report the average metrics of each level. For evaluation, we use three source views for every method to synthesize novel views. We report PSNR, SSIM and LPIPS to measure visual quality and depth MAE to measure the learned geometry quality. From Tab. 3, we observe that IBRNet outperforms in all three difficulty levels in terms of visual quality. Additionally, learned geometry quality isn’t highly correlated with the rendering visual quality in novel views. To sum up, our dataset provides great potential in learning category-level cross-scene NVS methods.
Figure 7: Relpose++ [34] pair-wise evaluation visualization. We show every image pair with its relative rotation predicted by Relpose.
Figure 8: Visualization of RGBD reconstruction surface from Neusfacto [83]. Original RGBD image samples are listed on the left and multi-view reconstructed surface is on the right for each example.
Eval. Type Categories Metrics #Frames
3 5 10 20
MST seen <<<15 deg. 57.4 55.1 51.4 47.4
<<<30 deg. 82.1 79.8 77.2 74.0
unseen <<<15 deg. 38.4 37.7 36.6 35.2
<<<30 deg. 62.5 61.8 60.4 59.0
Coord.Asc. seen <<<15 deg. 69.3 69.3 69.8 69.3
<<<30 deg. 85.3 85.3 85.4 85.3
unseen <<<15 deg. 46.0 46.2 46.9 46.5
<<<30 deg. 66.4 67.0 67.2 67.1
Sequential seen <<<15 deg. 51.9 45.1 36.0 26.9
<<<30 deg. 78.3 72.5 61.9 49.3
unseen <<<15 deg. 34.9 31.0 25.1 18.3
<<<30 deg. 59.1 54.4 46.4 37.1
Table 5: RelPose [84] camera evaluation results. We follow three evaluation types (MST, Coord.Asc., Sequential) propsed in [84] and report the average percent of rotation prediction errors in degrees both in training-seen categories and unseen ones.
Eval. Type Categories Metrics #Frames
2 3 5 8
Pairwise seen <<<15 deg. 69.6 68.3 67.3 66.6
<<<30 deg. 86.5 87.4 87.2 86.8
unseen <<<15 deg. 53.4 52.5 52.5 52.3
<<<30 deg. 74.1 74.5 75.3 75.4
Coord.Asc. seen <<<15 deg. 70.4 71.5 71.9 71.6
<<<30 deg. 86.7 87.9 88.5 88.7
unseen <<<15 deg. 52.9 55.4 54.9 54.8
<<<30 deg. 73.9 75.7 76.3 76.8
Cam.Center seen <<<0.2 $S$ 100.0 29.8 12.6 5.9
<<<0.3 $S$ 100.0 43.8 23.6 13.3
unseen <<<0.2 $S$ 100.0 30.5 12.4 5.6
<<<0.3 $S$ 100.0 44.3 23.2 12.5
Cam.Trans. seen <<<0.2 $S$ 22.3 11.9 4.8 2.2
<<<0.3 $S$ 30.1 20.4 10.8 6.2
unseen <<<0.2 $S$ 21.8 12.0 5.7 2.7
<<<0.3 $S$ 29.8 20.9 12.0 6.4
Table 6: RelPose++ [34] camera evaluation results. We follow four evaluation types (Pairwise, Coord.Asc., Cam.Center, Cam.Trans) proposed in [34] and report the average percent of rotation prediction errors in degrees both in training-seen categories and unseen ones. Notes: $S$ means scene scale defined in [34].
Depth Supervised NVS
To benchmark our dataset in NVS tasks, we also study the influences that depth supervision brings. We choose Instant-NGP [48] in single-scene NVS methods and both Pixel-NeRF [81] and MVSNeRF [6] in cross-scene NVS methods. Since depth supervision in NeRF-based methods provides more information on the scene, it should be beneficial for them to learn better representations. Our experiment results in Tab. 4 prove that. In our experiment setting, we add L1 depth loss to every algorithm and choose the best-performance depth loss weight for them. Compared with conducting NVS tasks only with RGB images, the performances of both Instant-NGP and Pixel-NeRF get boosted when training with depth loss. As for experiments of MVSNeRF, since we’ve already added depth information in the original training as the guidance in building rays, it turns out that the improvements when training with extra added depth loss is limited. In a nutshell, with depth priors, not only does single-scene NVS get improved due to more knowledge of the scene learned in training, but cross-scene NVS methods also learn better generalization capabilities. Depth images in our dataset have the potential to boost NVS more accurate and generalizable.
4.2 Camera Pose Estimation
In this section, we benchmark two data-driven methods RelPose [84] and RelPose++ [34] for inference of the relative camera poses from multi-view images in a sparse setting. Leveraging the given annotations of camera poses and large-scale category-level video in our dataset, we aim to learn generalizable viewpoint inference capability from training-seen categories to unseen ones. Since WildRGB-D dataset has a full and dense 360-degree camera trajectory (visualized in Figure. 2), we can provide both a large-scale database and various view settings to assist training. In our experiments, we divided totally 46 categories into training and testing categories. We also hold some videos in training categories for evaluation. We adopt evaluation settings described in [84, 34] and report results in Tab. 5 and Tab. 6. We observe that these two methods can generalize well to other scenarios both in known categories and unseen categories since the relative rotation estimation errors are in a reasonable range (also see Figure 7 for visualization). However, the error of translation prediction is comparatively large in RelPose++, which still poses challenges in this field. To sum up, with extensive category-level object scenes in our dataset, WildRGB-D dataset can serve as large-scale training sources for generalizable camera pose estimation algorithms to achieve remarkable results.
4.3 RGBD Object Surface Reconstruction
In our experiment setting of object surface reconstruction, algorithms need to utilize RGBD image sequence and central object masks to reconstruct the surface mesh of the central object. Reconstruction only using RGB images is also evaluated for comparison. For evaluation of reconstruction quality, we calculate the Chamfer Distance between the reconstruction mesh and aggregated object point cloud which is derived from object-masked depth images. Ten single object scenes are selected in each category for evaluations of Instant-NGP [48] and Neusfacto [83]. From the results in Tab. 7, we observe that with depth priors, the quality of reconstruction is better. Additionally, the performance of Neusfacto with RGBD is superior to Instant-NGP. Visualization of Neusfacto RGBD reconstructions are shown in Figure 8. In brief, our dataset provides an RGBD object reconstruction evaluation track, which might give rise to the development of more mature algorithms in this field.
RGB RGBD
Instant-NGP [48] Neusfacto [83] Instant-NGP [48] Neusfacto [83]
45.91/64.01 88.92/89.94 28.46/29.28 25.83/34.07
Table 7: RGBD object surface reconstruction results. Average of chamfer distance with standard deviation across selected categories are reported (Average/SD).
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Figure 9: Object 6D pose estimation visualization. We visualize the predicted category-level 6D pose on three categories in Wild6D [23] test set (bottle, bowl and mug) using models that only perform self-supervised training on the corresponding category of WildRGB-D Dataset. The ground truth bounding boxes are colored in green, and the predicted bounding boxes are colored in red.
[…]
2cm 5cm 2cm 5cm
Bottle Wild6D 93.2 85.2 71.3 79.4 79.8 90.9
ROW 93.3 70.9 34.1 48.8 47.9 78.8
Wild6D+ROW 93.3 85.8 71.9 78.6 81.7 91.7
Bowl Wild6D 98.3 90.4 66.1 70.0 86.8 94.6
ROW 98.3 91.8 33.8 35.7 86.3 93.5
Wild6D+ROW 98.4 91.8 40.3 42.0 87.5 93.7
Mug Wild6D 89.0 59.2 0.0 0.0 0.0 0.1
ROW 89.1 61.9 0.0 0.0 0.0 0.1
Wild6D+ROW 89.3 50.2 0.0 0.0 0.0 0.0
Table 8: Self-Supervised 6D Pose Estimation results. The evaluation results on Wild6D test dataset under three different settings in bottle, bowl and mug categories.
4.4 RGBD 6D Object Pose Estimation
We explore our dataset in self-supervised training for 6D pose estimation. We adopt the algorithm proposed in [85], which leverages category shape prior and learns by matching the correspondence between images and shapes. In our experiment, we evaluate the trained model on Wild6D [23] test set. Three different settings concerning training sets are adopted: 1) Wild6D training set; 2) WildRGB-D dataset; 3) Wild6D training set + WildRGB-D dataset. Common categories in these two datasets are selected for self-supervised training and evaluation. Results in Tab. 8 show that in the case of an out-of-distribution setting, where we trained only on our dataset and evaluated on a different dataset, some of the metrics are decreased in our experiments. This is mainly due to the different distribution of camera rotations in these two datasets (visualized in Fig. 4), where Wild6D doesn’t cover full 360 degrees and WildRGB-D dataset covers a larger pitch angle range in object 6D poses. However, we still witness some improvements in evaluations. Training with WildRGB-D dataset benefits in IOU evaluations and joint-dataset training improves rotation+translation evaluation in particular categories. Visualization of 6D pose estimation in Wild6D test set using models that only train in our dataset can be found in Figure 9, To summarize, our dataset provides large-scale category-level RGBD images sequences, serving as ample unsupervised training data, which has the potential to boost more accurate 6D pose estimation in the future.
5 Conclusion
The object-centric datasets in the computer vision community have mostly focused on RGB videos, while practical applications often involve depth as inputs or for better annotations. We collect the largest object-centric RGB-D video dataset WildRGB-D , where all videos are captured in cluttered scenes. It is composed of category-level RGB-D object videos taken using iPhones around the objects in 360 degrees, which contains around 8500 recorded objects and nearly 20000 RGB-D videos across 46 common object categories with three setups covering most scenarios. The dataset is well annotated with object masks, real-world scale camera poses, and reconstructed aggregated point clouds from RGBD videos. We set up four evaluation tracks with WildRGB-D, showing that the large-scale capture of RGB-D objects provides a large potential to advance 3D object learning. The current dataset does not come with annotations of the object 6D pose, which requires further crowd-sourcing effort. It will be one of our future efforts to collect this annotation for supervised training methods as well as evaluation. We are committed to releasing our dataset and evaluation code.

### v3

@@
 There are three types of videos: (i) Single object video where there is only one object presented on the table; (ii) Multi-object video where there are multiple objects presented at the same time; and (iii) Hand-object video where there is a static human hand grasping the object.
+More video types add variety, creating occlusion for objects in scenes, which are worthy study cases in some tasks.
 The collection of the WildRGB-D  dataset not only considers the cluttered background in the real world, but also the common scenarios where the objects are occluded by human hands.
@@
 ABO [14] none CAD mesh none 63 8k
-DTU [1] limited SfM mesh RGB N/A 124
-CO3D [53] full SfM pcl RGB 50 19k
-MVImgNet [82] limited SfM pcl RGB 238 220k
-Objectron [2] limited SfM pcl RGB 9 15k
+DTU [1] limited COLMAP mesh RGB N/A 124
+CO3D [54] full COLMAP pcl RGB 50 19k
+MVImgNet [84] limited COLMAP pcl RGB 238 220k
+Objectron [2] limited COLMAP pcl RGB 9 15k
 GSO [18] none scanner mesh none 17 1k
-OmniObject3D [73] limited scanner mesh RGB 190 6k
+OmniObject3D [75] full scanner mesh RGB 190 6k
 Choi et al. [13] limited sensor mesh* RGBD 9 2k
@@
 Depth Src.
-means where the depth information comes from, including CAD models, SfM COLMAP, scanner devices and depth sensor in iPhones.
+means where the depth information comes from, including CAD models, COLMAP, scanner devices and depth sensor in iPhones.
 pcl is the abbreviation of point cloud.
@@
 3D Object Datasets
-One representative kind of 3D object dataset is 3D synthetic dataset, like ShapeNet [5] and ModelNet40 [74], which consist of category-level objects.
+One representative kind of 3D object dataset is the 3D synthetic dataset, like ShapeNet [5] and ModelNet40 [76], which consist of category-level objects.
 3D-FUTURE [22] and ABO [14] datasets are typical of higher quality mesh with textures.
@@
 Objectron [2] provides rich annotations but only partial videos are fully 360 degree covered.
-CO3D [53] is a large-scale category-level dataset that annotates camera poses and depths with SfM tool COLMAP [56], which doesn’t provide depth information in real-world scale.
+CO3D [54] is a large-scale category-level dataset that annotates camera poses and depths with COLMAP [57], which doesn’t provide depths in real-world scale, and MVImgNet [84] is also a dataset similar to CO3D.
 Pascal3D [77] contains real-world 3D objects with pose annotations and CAD models in limited categories.
-Datasets collected with specialized hardware (scanner, dome, etc.) like GSO [18] and OmniObject3D [73] have more accurate 3D geometry models.
-However, they don’t have RGBD wild object videos collected.
+Datasets collected with specialized hardware (scanner, dome, etc.) like GSO [18] and OmniObject3D [75] have more accurate 3D geometry models and rendered depths from them.
+However, they don’t have RGBD wild object videos collected and lack real captured depths as well as background depths.
 In the aspect of RGBD object datasets, Wild6D [23] features RGBD image sequences and 6D pose annotations while lacking full 360 coverage and category types.
@@
 In order to generalize the NeRF representations to other scenes,  [83, 6, 70, 28, 43] learn latent 3D representations and priors from a bunch of existed scenes to help synthesize views across different scenes.
-Derived from original NeRF, [67, 79, 83, 15, 49, 72] leverage Sign Distance Function (SDF) and represent the 3D scene by implicit surface, which has a more clear object boundary definition.
+Derived from original NeRF, [68, 81, 85, 15, 50, 74, 72] leverage Sign Distance Function (SDF) and represent the 3D scene by implicit surface, which has a more clear object boundary definition.
+Recently, 3D Gaussian Splating [31] has become a competitive alternative to NeRF.
 WildRGB-D  dataset comprises various category-level objects and scenes on a large scale, which is suitable for novel view synthesis benchmarks and helps boost more mature reconstruction algorithms and generalizable 3D scene representations.
@@
 Data Collections
-In order to collect RGBD video on a large scale expediently and economically, we record with the help of iPhone front camera using Record3D  [^1] https://record3d.app/ App and rotate the camera around the object so that full 360-degree views of objects are captured with RGB images and the corresponding depth images.
-Camera rotating speed is controlled equably to ensure less blur in videos.
+In order to collect RGBD video on a large scale expediently and economically, we record with the help of an iPhone front camera using Record3D  [^1] https://record3d.app/ App and rotate the camera around the object so that full 360-degree views of objects are captured with RGB images and the corresponding depth images.
+Camera rotating speed is controlled equably by our collection setup to ensure less blur in videos.
 We select 46 common categories from well-known 2D datasets [17, 37].
@@
 Our WildRGB-D dataset has 3D annotations including camera poses in real-world scale, scene point clouds, and central object masks.
-In order to attain real-world scale camera poses, applying Structure-from-Motion (SfM) framework COLMAP [56] will result in relatively inaccurate SfM 3D camera pose annotations and the discrepancy of scaling in poses compared to real-world scaling, since SfM only leverages RGB image sequence and doesn’t take corresponding depths into account.
-The mature RGBD Simultaneous Localization and Mapping [19, 57] (SLAM) algorithms have the capability of exploiting the RGB images and depth information from depth sensor of mobile phones to reconstruct the 3D camera poses in real-world scale, which enables us to simply project the depth images and gain aggregated 3D point clouds (see Figure 5).
+In order to attain real-world scale camera poses, instead of relying on COLMAP [57] to first generate camera poses and then the depths using the poses, we generate more accurate camera poses with the mature RGBD Simultaneous Localization and Mapping [19, 58] (SLAM) algorithm, which leverages our captured depths.
+Additionally, it has the capability of exploiting the RGB images and depth information from the depth sensor of mobile phones to reconstruct the 3D camera poses in real-world scale, which is different from COLMAP depths, which are not in real-world scale.
+It enables us to simply project the depth images and gain aggregated 3D point clouds (see Figure 5).
 Then we manually check the quality of the aggregated 3D point cloud and exclude videos in which SLAM fails to get accurate camera poses.
@@
 NeRF-based methods perform better when we only concern with the recovery of central objects under object masks.
-What’s more, Mip-NeRF 360 performs best in learning single-scene geometry due to its special designs of proposal MLP and distortion loss.
-Visualization of single-scene NVS quality can be found in Figure 6.
-In brief, our dataset offers various and extensive categories and scenes for in-depth NVS experiments.
+What’s more, Mip-NeRF 360 performs best in learning single-scene geometry.
+Visualization can be found in Figure 6.
+In brief, our dataset offers extensive categories and scenes for in-depth NVS experiments.
 Cross-Scene NVS
-Apart from single-scene NeRF optimizations, we also evaluate Generalizable NeRFs: Pixel-NeRF [81], MVSNeRF [6] and IBRNet [69] in the cross-scene setting.
-For each category in our dataset, we select the same ten test scenes as single-scene NVS experiments and train in the remaining scenes of the same category to learn per-category latent representations.
+Apart from single-scene optimizations, we also evaluate Generalizable NeRFs: Pixel-NeRF [83], MVSNeRF [6] and IBRNet [70] in the cross-scene setting.
+For each category in our dataset, we select the same test scenes as single-scene NVS experiments and train in the remaining scenes of the same category to learn per-category latent representations.
 We divide the 46 categories into three difficulty levels and report the average metrics of each level.
-For evaluation, we use three source views for every method to synthesize novel views.
+For evaluation, we use three source views to synthesize novel views.
 We report PSNR, SSIM and LPIPS to measure visual quality and depth MAE to measure the learned geometry quality.
@@
 Figure 8: Visualization of RGBD reconstruction surface from Neusfacto [85]. Original RGBD image samples are listed on the left and multi-view reconstructed surface is on the right for each example.
+Depth Supervised NVS
+We also study the influences that depth supervision brings.
+We choose Instant-NGP [49] in single-scene NVS methods and both Pixel-NeRF [83] and MVSNeRF [6] in cross-scene NVS methods.
+Our experiment results in Tab.
+4 prove that depth supervision is beneficial for these methods to learn better representations.
+In our experiment setting, we add L1 depth loss to every algorithm and choose the best-performance depth loss weight for them.
+Compared with conducting NVS tasks without depths, the performances of both Instant-NGP and Pixel-NeRF get boosted when training with depth loss.
+As for experiments of MVSNeRF, since we’ve already added depth information in the original training as the guidance in building rays, it turns out that the improvements when training with extra added depth loss are limited.
+In a nutshell, with depth priors, both single-scene NVS and cross-scene NVS methods learn better generalization capabilities, boosting NVS more accurate and generalizable.
+4.2 Camera Pose Estimation
+In this section, we benchmark two data-driven methods RelPose [86] and RelPose++ [35] for inference of the relative camera poses from multi-view images in a sparse setting.
+Leveraging the given annotations of camera poses and large-scale category-level video in our dataset, we aim to learn generalizable viewpoint inference capability from training-seen categories to unseen ones.
+Since WildRGB-D dataset has a full and dense 360-degree camera trajectory, we can provide both a large-scale database and various view settings to assist training.
+In our experiments, we divided totally 46 categories into training and testing categories.
+We also hold some videos in training categories for evaluation.
+We adopt evaluation settings described in [86, 35] and report results in Tab.
+5 and Tab.
+6.
+We observe that these two methods can generalize well to other scenarios both in known categories and unseen categories since the relative rotation estimation errors are in a reasonable range (also see Figure 7 for visualization).
+However, the error of translation prediction is comparatively large in RelPose++, which still poses challenges in this field.
+To sum up, WildRGB-D dataset can serve as large-scale training sources for generalizable camera pose estimation algorithms to achieve remarkable results.
+4.3 RGBD Object Surface Reconstruction
+In our experiment setting of object surface reconstruction, algorithms need to utilize RGBD image sequence and central object masks to reconstruct the surface mesh of the central object.
+Reconstruction without depths is also evaluated for comparison.
+For evaluation of reconstruction quality, we calculate the Chamfer Distance between the reconstruction mesh and aggregated object point cloud which is derived from object-masked depth images.
+Ten single object scenes are selected in each category for evaluations of Instant-NGP [49] and Neusfacto [85].
+From the results in Tab.
+7, we observe that reconstruction is better with depth priors.
+Additionally, the performance of Neusfacto with RGBD is superior to Instant-NGP, which shows that depths help the sdf-based method Neusfacto learn the correct object boundary and boost the performance more compared with Instant-NGP.
+The average deviation is high due to the varied reconstruction qualities across different categories in the dataset.
+Visualization of Neusfacto RGBD reconstructions are shown in Figure 8.
+In brief, our dataset provides an RGBD object reconstruction evaluation track, boosting the development of more mature algorithms in this field.
+[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Figure 9: Object 6D pose estimation visualization.
+We visualize the predicted category-level 6D pose on three categories in Wild6D [23] test set (bottle, bowl, and mug) using models that only perform self-supervised training on the corresponding category of WildRGB-D Dataset.
+The ground truth bounding boxes are colored in green, and the predicted bounding boxes are in red.
 Eval. Type Categories Metrics #Frames
@@
 Notes: $S$ means scene scale defined in [35].
-Depth Supervised NVS
-To benchmark our dataset in NVS tasks, we also study the influences that depth supervision brings.
-We choose Instant-NGP [48] in single-scene NVS methods and both Pixel-NeRF [81] and MVSNeRF [6] in cross-scene NVS methods.
-Since depth supervision in NeRF-based methods provides more information on the scene, it should be beneficial for them to learn better representations.
-Our experiment results in Tab.
-4 prove that.
-In our experiment setting, we add L1 depth loss to every algorithm and choose the best-performance depth loss weight for them.
-Compared with conducting NVS tasks only with RGB images, the performances of both Instant-NGP and Pixel-NeRF get boosted when training with depth loss.
-As for experiments of MVSNeRF, since we’ve already added depth information in the original training as the guidance in building rays, it turns out that the improvements when training with extra added depth loss is limited.
-In a nutshell, with depth priors, not only does single-scene NVS get improved due to more knowledge of the scene learned in training, but cross-scene NVS methods also learn better generalization capabilities.
-Depth images in our dataset have the potential to boost NVS more accurate and generalizable.
-4.2 Camera Pose Estimation
-In this section, we benchmark two data-driven methods RelPose [84] and RelPose++ [34] for inference of the relative camera poses from multi-view images in a sparse setting.
-Leveraging the given annotations of camera poses and large-scale category-level video in our dataset, we aim to learn generalizable viewpoint inference capability from training-seen categories to unseen ones.
-Since WildRGB-D dataset has a full and dense 360-degree camera trajectory (visualized in Figure.
-2), we can provide both a large-scale database and various view settings to assist training.
-In our experiments, we divided totally 46 categories into training and testing categories.
-We also hold some videos in training categories for evaluation.
-We adopt evaluation settings described in [84, 34] and report results in Tab.
-5 and Tab.
-6.
-We observe that these two methods can generalize well to other scenarios both in known categories and unseen categories since the relative rotation estimation errors are in a reasonable range (also see Figure 7 for visualization).
-However, the error of translation prediction is comparatively large in RelPose++, which still poses challenges in this field.
-To sum up, with extensive category-level object scenes in our dataset, WildRGB-D dataset can serve as large-scale training sources for generalizable camera pose estimation algorithms to achieve remarkable results.
-4.3 RGBD Object Surface Reconstruction
-In our experiment setting of object surface reconstruction, algorithms need to utilize RGBD image sequence and central object masks to reconstruct the surface mesh of the central object.
-Reconstruction only using RGB images is also evaluated for comparison.
-For evaluation of reconstruction quality, we calculate the Chamfer Distance between the reconstruction mesh and aggregated object point cloud which is derived from object-masked depth images.
-Ten single object scenes are selected in each category for evaluations of Instant-NGP [48] and Neusfacto [83].
-From the results in Tab.
-7, we observe that with depth priors, the quality of reconstruction is better.
-Additionally, the performance of Neusfacto with RGBD is superior to Instant-NGP.
-Visualization of Neusfacto RGBD reconstructions are shown in Figure 8.
-In brief, our dataset provides an RGBD object reconstruction evaluation track, which might give rise to the development of more mature algorithms in this field.
 RGB RGBD
@@
 Table 7: RGBD object surface reconstruction results. Average of chamfer distance with standard deviation across selected categories are reported (Average/SD).
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Figure 9: Object 6D pose estimation visualization.
-We visualize the predicted category-level 6D pose on three categories in Wild6D [23] test set (bottle, bowl and mug) using models that only perform self-supervised training on the corresponding category of WildRGB-D Dataset.
-The ground truth bounding boxes are colored in green, and the predicted bounding boxes are colored in red.
-[…]
+Category Datasets IOU0.25 IOU0.5 5 deg. 5 deg. 10 deg. 10 deg.
 2cm 5cm 2cm 5cm
@@
 We are committed to releasing our dataset and evaluation code.
+Acknowledgment
+This project was supported, in part, by the Amazon Research Award, the Qualcomm Innovation Fellowship, the Intel Rising Star Faculty Award, and the CISCO Faculty Award.